OpenStudy (anonymous):

help please? simplify (x - 1)(x2 + 10x + 4) x3 + 10x2 + 4x - 1 x3 + 9x2 - 6x - 4 x3 + 10x2 + 4x - 4 x3 + 9x2 - 6x - 1

OpenStudy (apoorvk):

(x-1)(x^2+10x+4) = x(x^2 + 10x + 4) -1(x^2 + 10x + 4) = x^3 + 10x^2 + 4x -x^2 -10x - 4 = x^3 + 9x^2 -6x -4 ....{ans}